OpenStudy (anonymous):

The sum of two numbers is 34. The smaller number is 12 less than the larger number. What are the numbers?

OpenStudy (anonymous):

let x and y be numbers ,x>y x+y=34 x-y=12 add 2x=?

OpenStudy (anonymous):

hi welcome to openstudy @neyney1997 :) let's do it step by step. let x be our big number and y be our smaller number. "sum of two numbers is 34" so like what surjithayer said, x + y = 34 let this be our equation 1. "smaller number is 12 less than the larger number" so y = x-12 or it will be the same as x-y = 12 let this be our equation 2. Now, solve the values of x and y using elimination method. We will eliminate one variable to solve for another variable by adding equation 1 and equation 2. Can you solve this part: